### v1

# Towards Contextual Sensitive Data Detection

Abstract
The emergence of open data portals necessitates more attention to protecting sensitive data before datasets get published and exchanged. While an abundance of methods for suppressing sensitive data exist, the conceptualization of sensitive data and methods to detect it, focus particularly on personal data that, if disclosed, may be harmful or violate privacy. We observe the need for refining and broadening our definitions of sensitive data, and argue that the sensitivity of data depends on its context. Based on this definition, we introduce two mechanisms for contextual sensitive data detection that consider the broader context of a dataset at hand. First, we introduce type contextualization, which first detects the semantic type of particular data values, then considers the overall context of the data values within the dataset or document. Second, we introduce domain contextualization which determines sensitivity of a given dataset in the broader context based on the retrieval of relevant rules from documents that specify data sensitivity (e.g., data topic and geographic origin). Experiments with these mechanisms, assisted by large language models (LLMs), confirm that: 1) type-contextualization significantly reduces the number of false positives for type-based sensitive data detection and reaches a recall of 94% compared to 63% with commercial tools, and 2) domain-contextualization leveraging sensitivity rule retrieval is effective for context-grounded sensitive data detection in non-standard data domains such as humanitarian datasets. Evaluation with humanitarian data experts also reveals that context-grounded LLM explanations provide useful guidance in manual data auditing processes, improving consistency. We open-source mechanisms and annotated datasets for contextual sensitive data detection at https://github.com/trl-lab/sensitive-data-detection.
1 Introduction
Data sharing underpins reuse across sciences, enterprises, and governments through public and private portals [borgman2025data, worth2024datatransparency, zhang2024datamarketssurvey, wilkinson2016fair, brickley2019googledataset]. With the rise of open data sharing portals, comes the need to protect sensitive data from undesired and harmful use. Sensitive data includes direct personal identifiers (e.g., names, emails [gdpr]) as well as indirect or contextual information that becomes a potential risk under certain conditions [unocha2025]. For example, facility locations or military logs may be sensitive depending on geopolitics or timing. This illustrates the complexity of sensitive data detection and the need for nuanced, context-dependent methods. Beyond operational and ethical concerns of publishing sensitive data, Large Language Models (LLMs) increase the risks of leakage: trained on web data, they may memorize and reproduce sensitive information [subramani2023piihuggingface, worth2024datatransparency, carlini2021extracting, lukas2023analyzing]. Studies show sensitive data is often underspecified in documentation on open data portals like HuggingFace [yang2024navigatingdatasetdocumentationsai, croissant2024benjelloun, worth2024datatransparency], illustrating the importance of accurate methods and tools for (semi-)automated sensitive data protection, particularly in open data portals.
Most work on sensitive data detection methods, however, focuses on detecting person-level data, particularly personal identifiable information (PII) [KUZINA2023119924, subramani2023piihuggingface, croissant2024benjelloun]. For unstructured data, PII detection relies on named entity recognition [subramani2023piihuggingface]; similar approaches exist for tabular data [raman2001potter, Hulsebos2019Sherlock], which is the focus in this paper. Yet sensitivity extends far beyond explicit PII. kober2023whatissensitive propose a taxonomy of always-, combination-, context-, and value-sensitive data. For instance, an address may be harmless in the context of a public organizational website but expose a private address in another document, causing naïve type-based detection yield false positives. Likewise, hospital geo-coordinates may be harmless in one region but sensitive in conflict zones [unocha2025], calling for contextualization beyond detecting personal-sensitive data. These limitations call for a more nuanced perspective and framework for assessing sensitive data. LLMs offer new potential for sensitivity detection by reasoning over structured data. Studies show they can interpret tables with carefully designed prompts [fang2024largelanguagemodelsllmstabular], achieving strong results in tasks such as text-to-SQL and table retrieval [fang2024largelanguagemodelsllmstabular, ji2025target]. These capabilities suggest that LLMs can leverage schema, values, and external knowledge to interpret tabular dataset semantics [Hulsebos2019Sherlock].
In this paper, we first revisit the concept of sensitive data, arguing for a more contextual perspective. We introduce two new LLM-assisted mechanisms for contextual sensitive data detection. First, we introduce the detect-then-reflect mechanism for type contextualization of data values within the entire document or tabular dataset to assess their sensitivity. Experiments with real-world datasets containing personal identifiable information show that this mechanism yields a strong improvement in precision while maintaining high recall. Second, the retrieve-then-detect mechanism for domain contextualization which grounds the assessment in domain-specific sensitivity specifications. Evaluations illustrate the potential of this mechanism, achieving high recall and improved consistency for detecting non-personal sensitive data. Importantly, the context-grounded domain-specific explanations were deemed highly useful by experts in humanitarian data assessment in a case-study.
2 Background
Conceptualization of Sensitive Data
The notion of “sensitive data” is multifaceted, spanning legal, theoretical, and domain-specific views. Legal frameworks often equate sensitivity with personal identifiable information. The GDPR defines personal data broadly as “any information relating to an identified or identifiable natural person,” specifying sensitive categories like health, ethnicity, or politics requiring extra protection [gdpr]. Similarly, HIPAA defines identifiers that make health information “Protected Health Information (PHI)” when linked to individuals [hipaa]. These protocols provide a baseline: data that directly or indirectly identifies a person is sensitive. Yet static, type-level definitions are insufficient for modern governance [Quinn_Malgieri_2021], as they emphasize direct identifiers while overlooking non-personal but risky data such as geolocation data, community-level attributes, or institutional records. Sensitivity is often contextual: seemingly harmless data can become sensitive when combined, repurposed, or shared across contexts [malkin2022contextualintegrity].
Theoretical perspectives deepen this nuance. Nissenbaum’s Contextual Integrity frames sensitivity as the appropriateness of information flows within social norms [nissenbaum2004privacy]. Violations occur when flows breach expectations, not from the data type itself. Building on this, researchers advocate context-grounded definitions [malkin2022contextualintegrity, shvartzshnaider2019going]. For instance, shopping data may be ordinary in retail but sensitive in behavioral profiling. Recent taxonomies further refine this view. kober2023whatissensitive distinguish: (1) always-sensitive data (e.g., direct identifiers); (2) combination-sensitive (sensitive when linked together); (3) context-sensitive (sensitivity by use); and (4) value-sensitive (sensitivity from specific values). This shows that not all personal is equally sensitive, and not all sensitive data is on individuals.
In short, conceptualizations have shifted from static identifiers to context-dependent models, where frameworks like Contextual Integrity and multi-layer taxonomies emphasize context, use, and value. This motivates our focus: effective data protection must ensure coverage beyond personal-level information toward context-aware, domain-specific approaches.
Methods for Detecting Sensitive Data
Most automated sensitive data detection methods concentrate on detecting sensitive data of individuals. Widely adopted industry tools like Microsoft Presidio [presidio] and Google Cloud DLP [google_DLP] are largely rule-based, relying on regular expressions, keyword lists, and simple machine learning heuristics to recognize well-defined patterns like email addresses or phone numbers. Research has advanced towards more semantic understanding. Hulsebos2019Sherlock introduced Sherlock, a deep learning model that classifies types of table columns, including numerous PII, by learning from column-level representations. KUZINA2023119924 develop this further with CASSED, a BERT-based model that integrates column headers, data types, and sample values to classify columns into sensitive categories. These approaches outperform regular expressions and dictionary-based baselines, especially on noisy or inconsistently formatted data. Recently, Large Language Models (LLMs) have emerged as powerful tools for sensitive data detection, as particularly shown for PII detection outperforming both rule-based and traditional machine learning methods in accuracy and adaptability[shen2025enhancingdeidentificationpersonallyidentifiable, Yang2023Exploring]. However, a key limitation remains: LLMs excel at identifying known PII types but struggle with assessing contextual sensitivity, such as domain-specific sensitive types, unless explicitly provided with the relevant operational context.
We identify a significant gap in detecting sensitive data: methods focus on detecting personal sensitive data, overlooking non-personal data that is sensitive due to its context. Our work bridges this gap by developing a framework that addresses both personal and non-personal sensitive data detection through contextualization.
3 Revisiting the Concept of Sensitive Data
The main conceptual, legal, and methodological focus regarding sensitive data is on personal- and individual level data, including personal identifiable information (PII) which refers to data that can identify an individual [gdpr]. Regulatory data protocols, such as GDPR [gdpr] and HIPAA [hipaa], specify standards for data sensitivity, hence provide concrete guidelines for data protection measures. However, we argue for the need of a more contextual view of sensitive data (Def. 1) for developing new detection methods.
Definition 1(Contextual Sensitive Data)
Data whose sensitivity depends on external factors and requires protection due to how, by whom, and in what context it can be misused.
Advancing this definition, we posit two contextual aspects that matter in order to assess sensitivity of data: type contextualization and domain contextualization.
Following the existence of generic standards and protocols, sensitive data detection methods often focus on detecting data types reflecting personal sensitive data (e.g., names, addresses, phone numbers, birth dates). But the appearance of a name or address does not immediately make a data value sensitive. For example, an address may reflect the address of a person or organization, each having significantly different consequences, hence different sensitivity levels. We term this type contextualization: whether a data value corresponding to a sensitivity type of interest (e.g. name) truly is sensitive requires considering the context in which it appears in the entire dataset (e.g., the entire document or table).
Besides a more fine-grained contextual notion of a data value its sensitivity, we also argue for a more coarse-grained contextual perspective. Nissenbaum introduced the theory of Contextual Integrity [nissenbaum2004privacy], which states that privacy risks arise when data is used outside its intended setting, particularly addressing personal privacy. This contextual perspective extends beyond personal privacy, in domains such as humanitarian aid, national security, and company-sensitive context. We pose that data protection should account for risks beyond identity exposure, for example, harming vulnerable populations or organizations. We refer to this broader perspective on data sensitivity as domain contextualization: sensitivity emerges from the dataset’s nature and potential misuse.
4 Toward Contextual Sensitive Data Detection
Protecting sensitive data generally follows a three-step procedure: ➀ specifying sensitive data, ➁ detecting sensitive data, ➂ remedying data sensitivity (Figure 1). In this work, we focus on the second step, which has received little attention in research. Existing methods for detecting sensitive data suffer from two key limitations to address the contextual nature of sensitive data:
  1. 1.
    Insufficient specificity: sensitive data detection methods treat all values of types of interest as equally sensitive, while the subject of the value that implies its sensitivity. The lack of specificity leads to low precision hence high false positive rates.
  2. 2.
    Insufficient coverage: sensitive data detection methods typically focus on personal-level data, particularly PII, and do not generalize to domain-specific sensitive data. This lack of coverage leads to low recall of methods, hence leaving sensitive data unprotected.
Figure 1: Data protection process, starting with the specification of sensitivity, detection sensitive data in inputs accordingly, followed by remedying sensitive data by suppression methods.
The contextualization framework introduced in Section 3 addresses these limitations. Type contextualization would improve specificity as potential sensitive values are only considered sensitive based on the broader context of the value. In turn, domain contextualization ensures higher coverage of sensitive data assessment by grounding the assessment in domain-specific context. In this paper, we introduce two concrete mechanisms to address these contextual aspects of sensitive data: the detect-then-reflect mechanism for type contextualization and the retrieve-then-detect mechanism for domain contextualization. In what follows, we specifically focus on tabular data but these mechanisms extend to other modalities, such as free-form text documents.
(a) The detect-then-reflect mechanism for detecting sensitive table columns using type contextualization to increase the precision. This mechanism entails: ➀ detecting potentially sensitive columns based on their type, ➁ having an LLM reflect on the actual sensitivity of the column within the context of the entire table.
(b) The retrieve-then-detect mechanism for detecting sensitive data that requires domain contextualization. This mechanism entails: ➀ retrieving relevant contextual information such as rule-based heuristics, ➁ using an LLM to detect sensitive data using its parametric knowledge and the retrieved expert heuristics.
Figure 2: Mechanisms for detecting sensitive data in a contextual manner.
4.1 Type contextualization: detect-then-reflect
Existing methods typically classify sensitive data based on signals indicating a type of interest, e.g. names, addresses, or phone numbers. If a value matches this type, it is flagged as sensitive. We introduce a two-step pipeline for detecting type-informed sensitive data based on a detect-then-reflect procedure (Figure 2(a)). In the initial detection step, potentially sensitive columns are identified by first classifying columns corresponding to types of interest that are specified upfront, e.g. email, phone number, based on column names and data values. This step prioritizes recall in order to avoid missing any columns that may carry data of the type that can potentially be sensitive. In the second step, reflection, the model assesses the columns in the context of the full table. We prompt an LLM with the table, including the column’s detected type. The LLM determines whether the column is truly sensitive, taking into account inter-column relationships and data across the entire table. This step is crucial for mitigating false positives from the detection stage by incorporating inter-column relationships and table semantics.
4.2 Domain contextualization: retrieve-then-detect
While type-based data sensitivity can often be inferred from structural context, the data sensitivity may depend on domain-specific and geopolitical factors that lie beyond the data itself. For such cases, requiring domain contextualization, we introduce the retrieve-then-detect mechanism (Figure 2(b)). This mechanism considers external contextual information, such as rule-based heuristics specified in data governance documents, in the classification process. Relevant documents are retrieved based on, for example, the origin of a dataset and included in the prompt alongside the table and column under analysis. The LLM leverages both its internal parametric knowledge and the retrieved external data sensitivity instructions to assess whether a column contains data that should be considered sensitive in the given operational context. Unlike type-informed detection, where there is a predefined list of entities (e.g. email, phone number) as relevant for for example PII detection, domain-contextual sensitive data has no fixed ontology. This mechanism establishes domain-specific column sensitivity based on the broader table context and retrieved contextual knowledge.
5 Experimental Setup
Here, we outline the implementation details, the LLMs used, baselines, datasets, and metrics to evaluate our proposed framework for contextually sensitive data detection. Code and datasets can be found at https://github.com/trl-lab/sensitive-data-detection.
5.1 Method implementation
5.1.1 Type-contextualization
In the detection step, given the well-defined nature of PII types, we parse the problem as a multiclass classification task over a taxonomy of personal identifiable types (e.g., email_address, generic_identifier) derived from GDPR [gdpr] and UN OCHA guidelines [OCHA2021Guidelines] (see the code repository for a complete list). The LLM is prompted with a column name and five values, to return only the most appropriate type or None. We consider zero-shot and fine-tuned approaches. The zero-shot method leverages the model’s inherent ability to follow instructions, while the fine-tuned approach enhances performance by training the models on a synthetic dataset of 1,000 labeled columns. Fine-tuning is done using Low-Rank Adaptation [LORA21_HU] for parameter efficiency. The models are optimized with AdamW for a single epoch using an 80/20 train-validation split.
In the reflection phase, the LLM re-evaluates each candidate column corresponding to types of interest within the context of the entire table. We provide the LLM with a markdown-formatted representation of the full table, including headers and five sample rows per column, along with the specific type detected for the target column and, if applicable, the entities of other columns which are detected as PII entities. The LLM is then instructed to assign one of three contextual sensitivity levels: non_sensitive, referring to information that cannot identify a person (e.g., aggregate data, organization names); moderate_sensitive, referring to data that could potentially identify a person when combined with other attributes (e.g., demographics, partial information); and high_sensitive, referring to data that definitively identifies a person (e.g., full name, email, national ID).
5.1.2 Domain-contextualization
For the retrieval step, we aim to identify relevant heuristics indicating sensitivity rules from domain-specific data governance documents in order to integrate this external domain knowledge in the LLM’s context. In our experiments we focus on a case study on the Humanitarian Data Exchange (HDX) [hdx], hence on humanitarian datasets. The system retrieves relevant policy fragments from Information Sharing Protocols (ISPs) based on the dataset’s geographical context (e.g., country). We use GPT-4o-mini to extract the specific sensitivity rules from these documents (e.g., data types for low, moderate, high, severe sensitivity), which we store in a dictionary. If no country-specific ISP is identified, a default fallback template based on cross-country consensus is used.
In the detection phase, the LLM is provided with a curated prompt containing table context (headers and five representative rows), relevant ISP extracts containing explicit instructions to classify each column’s sensitivity according to the ISP guidelines while justifying its decision by directly citing the relevant policy passages. Based on the policy definitions, the model assigns one of four sensitivity levels, non_sensitive, moderate_sensitive, high_sensitive, or severe_sensitive. This mechanism ensures that sensitivity assessments are grounded in explicit domain policies, thereby enhancing both interpretability and practical usability for domain experts.
5.2 Model Selection
We evaluate closed-source and open-source models, to understand trade-offs between these options. From closed-source options, we select the GPT-4o-mini [openai2024gpt4ocard] model for its good performance and cost trade-off. We focus on open-source LLMs that enable local deployment, for which we select the Gemma family (2 and 3) [gemma_2025, Gemma2Paper] and Qwen3 in 8B and 14B sizes [yang2025qwen3technicalreport]. Qwen3 14B was selected given the strong performance of its predecessor Qwen2.5’s exceptional performance on benchmarks such as MMLU Pro  [^1] https://huggingface.co/spaces/TIGER-Lab/MMLU-Pro, demonstrating strong reasoning capabilities and robust general understanding. Furthermore, Gemma2 9B and Gemma3 12B were included to strike a balance between computational efficiency and performance, and their support for quantized deployment. We also include versions of Qwen3 8B and Gemma2 9B that are fFgine-tuned on a synthetic dataset of 1K table columns hand-labeled with 27 PII types, indicated with “FT” suffix.
5.3 Baselines
We compare our methods with two common baselines for sensitive data detection. We include Google Cloud DLP [google_DLP], a rule-based system that combines built-in infoTypes, predefined detectors for common PII entities, which are associated custom-defined dictionaries and regular expressions. For our experiments, we configure the minimum likelihood threshold to “Likely”. The second baseline is Microsoft Presidio, an open-source tool that leverages regular expressions, pattern-based rules, context windows, and ML-based named entity recognition to detect PII types. To enable consistent evaluation, each Presidio type has been mapped to a corresponding category in our unified taxonomy (this mapping can be found in the code repository). Both systems output column-level predictions and are widely adopted in enterprise settings. However, neither tool supports contextual reflection or table-wide reasoning; any match to a known PII pattern is automatically treated as sensitive.
5.4 Datasets
To evaluate the contextual sensitive data detection mechanisms we use a combination of synthetic and real-world datasets, to maintain ground-truth validity for sensitive data detection in generic context and the domain-specific context of humanitarian datasets.
  • Type contextualization with sensitive PII: We extracted 66 anonymized tables (2,061 columns) from the GitTables corpus [hulsebos2023gittables] representing a variety of real-world database schemas. Each column was manually annotated with (i) its PII entity type and (ii) its contextual sensitivity level (non-sensitive, moderate or highly sensitive).
  • Domain contextualization with humanitarian tables: To evaluate the retrieve-then-detect mechanism, we generated 9 synthetic tables with sensitive data closely resembling typical humanitarian data (e.g., for displacement tracking, public health monitoring) for different countries. Each table contains columns that could be potentially sensitive within country-specific Information Sharing Protocols (ISPs). Humanitarian data quality experts labeled each column with a sensitivity level (low, moderate, high, severe). We also add 14 real tables from the Humanitarian Data Exchange [^2] https://data.humdata.org/ that do not contain any sensitive data, to evaluate the false positive rate.
5.5 Metrics
We evaluate along three axes: PII detection, contextual reflection for PII, and non-PII sensitivity detection. To account for class imbalance, we report both weighted (support-weighted) and macro averages. Weighted averages reflect overall performance across examples, while macro averages treat each class equally and reduce dominance by frequent classes. Unless otherwise noted, we focus on weighted averages.
Performance is assessed at both the column and table level. At the column level (type contextualization), PII detection is treated as a multiclass task over 27 PII types (plus a “None” category), while sensitivity reflection reduces to a binary decision between sensitive (moderate/high) and non-sensitive (low). At the table level (domain contextualization), a table is classified as sensitive if it contains any moderate/high/severe sensitive information, resulting in a binary task (sensitive vs. not). Precision, recall, and F1 are used as evaluation metrics.
For labeling, sensitivity levels are binarized: low is treated as non-sensitive, while moderate, high, and severe are grouped as sensitive. This aligns with operational practice where any moderate or higher sensitivity triggers safeguards, consistent with UN OCHA guidance that such data warrants protective measures [unocha2025].
6 Results of contextual sensitive data detection
We present a comprehensive evaluation of our proposed framework for contextually sensitive data detection in tabular datasets, through type- and domain-contextualization. Type Contextualization experiments instrumented by the detect-then-reflect mechanism are focused on personal-identifiable information (PII) and Domain Contextualization instrumented by the retrieve-then-reflect mechanism which focus on less standard data semantics through a humanitarian case-study.
6.1 Type contextualization improves specificity.
First, we assess the detect-then-reflect mechanism for detecting sensitive data through type contextualization. We report on baselines representing existing state-of-the-art tools for detecting sensitive data which automatically treat all columns of interest types as sensitive above the line. Below the line, we report the contextual LLM-assisted method that detects columns of the types of interest which are directly labeled as sensitive without reflection (left column), and are labeled as sensitive only if this is the output of the reflection step (right column).
Table 1: Type contextualization for PII sensitive data detection, with baselines and the detect-then-reflect mechanism. The LLMs are shown with and without the reflection step. We observe that reflection consistently improves precision and F1.
System / Model Without Reflection With Reflection
Prec. Rec. F1 Prec. Rec. F1
Google DLP 0.531 0.628 0.576
Presidio 0.520 0.618 0.565
GPT-4o-mini 0.856 0.639 0.732 0.938 0.632 0.755
Gemma 2 9B 0.740 0.819 0.778 0.800 0.792 0.796
Gemma 3 12B 0.487 0.941 0.642 0.753 0.806 0.779
Qwen3 8B 0.742 0.868 0.800 0.749 0.868 0.804
Qwen3 14B 0.565 0.972 0.714 0.732 0.941 0.824
Qwen3 8B FT $\rightarrow$ GPT-4o-mini 0.902 0.861 0.881
Without reflection, LLMs mark any detected PII-type column as sensitive, leading to high recall but lower precision; Gemma 3 12B and Qwen3 14B reach a recall of 0.94 and 0.97 but precision of only 0.49 and 0.57, respectively. GPT-4o-mini achieves higher precision, 0.86, at lower recall of 0.64. With reflection using full-table context, substantially improves precision across models, e.g., Qwen3 14B increases to 0.73 precision with 0.94 recall, trading a small decrease in recall for a significant gain on precision. A modular pipeline combining Qwen3 8B for type detection with GPT-4o-mini for reflection achieves F1 0.88, but is closely matched by Qwen3 14B which yields higher recall (0.94 compared to 0.86). These results illustrate the stark difference compared to existing baselines reported in the upper left quadrant, which suffer from insufficient recall ánd precision. Comparing the detect-then-reflect mechanism powered by open-source LLMs with commonly used pattern-matching tools DLP and Presidio, we observe a significant improvement in both precision and recall, addressing the key limitations of existing tools for deployment in practice. Overall, the detect-then-reflect mechanism corrects the naive assumption that all columns with potentially sensitive types are sensitive, boosting precision while maintaining recall.
6.1.1 Ablation: reflection without detection.
To investigate the value of the type-detection step, we also task models with identifying sensitive columns using directly the full table context without first detecting the presence of columns of specific types (Table 1). The results in Table 2 illustrate that LLMs yield good performance for sensitive columns using only table context without prior type detection, but generally yield significant improvement across models when explicit types are detected first. Qwen3 14B achieves the highest F1 (0.82) followed by Gemma 2 9B (F1 0.81). GPT-4o-mini excels in precision (0.96) but has low coverage with a recall of 0.56. Comparing the performance with the detect-then-reflect pipeline in Table 1, we observe that the reflection-only approach, without type detection, yields particularly a lower coverage, therefore leaving more sensitive data undetected. For example, looking closer into the performance of the best performing models, the recall of Qwen3 14B drops from 0.94 to 0.89, and Gemma2 9B from 0.79 to 0.75. These results indicate that LLMs can reason about sensitivity from the table alone, but the full detect-then-reflect mechanism yields preferable trade-off in precision and recall.
Table 2: Reflection-only performance without prior type detection step, for identifying contextually sensitive columns. Evaluated across models using precision, recall, and F1 scores.
Model Precision Recall F1
GPT-4o-mini 0.964 0.562 0.711
Gemma 2 9B 0.878 0.750 0.809
Qwen3 8B 0.735 0.781 0.758
Qwen3 14B 0.761 0.885 0.819
6.2 Domain contextualization provides coverage and context-grounded explainability.
Shifting towards contextual sensitive data which cannot be assessed through a fixed set of types, we evaluate the retrieve-then-detect mechanism on a combined set of synthetic and real humanitarian tables from the HDX platform. Without external domain knowledge, i.e. sensitivity heuristics obtained through context retrieval, all models exhibit a highly conservative bias, achieving perfect recall but all with precision lower than 0.56 (Table 3). Integrating retrieved domain-specific context, in this case we leverage information sharing protocols (ISP’s) for humanitarian data sharing, into the prompt consistently improves precision. GPT-4o-mini’s precision increases from 0.47 to 0.69 (F1=0.82), and Qwen3 14B’s from 0.56 to 0.64 (F1=0.78), both maintaining perfect recall (1).
Table 3: Table-level classification performance with and without context-grounding in humanitarian data sensitivity rules. The retrieved rules improve precision while maintaining high recall.
System / Model No Domain Knowledge With Domain Knowledge
Prec. Rec. F1 Prec. Rec. F1
All-tables-sensitive (baseline) 0.375 1.000 0.545
GPT-4o-mini 0.474 1.000 0.643 0.692 1.000 0.818
Gemma 2 9B 0.375 1.000 0.545 0.429 1.000 0.600
Gemma 3 12B 0.529 1.000 0.692 0.500 1.000 0.667
Qwen3 8B 0.562 1.000 0.720 0.778 0.778 0.778
Qwen3 14B 0.562 1.000 0.720 0.643 1.000 0.783
Qualitative analysis reveals that the retrieval mechanism enables models to produce interpretable justifications, thereby enhancing the transparency and auditability of the sensitivity assessments. The below excerpt shows a context-grounded LLM assessment on a table titled “Reported location on NFIs suppliers” with a description of the data at hand (1), along with the LLM’s sensitivity assessment (2), and explanation grounded in retrieved ISP rules (3).
1. The column ‘‘nfi_supplier_loc- value -from_Iran’’ indicates the origin of NFIs suppliers.
2. This information, even with country-level disaggregation, doesn’t reveal personal or sensitive data. It relates to the location of suppliers, which is generally not considered sensitive.
3. It aligns with the ISP’s guideline on ‘‘Facility data (e.g., health, education, water points) at national or regional level, unless explicitly restricted by clusters.’’ Since this column is about locations of suppliers and not about individuals or specific operations, it is categorized as non-sensitive.
The model outputs containing sensitivity classifications and explanations were also reviewed by two humanitarian data quality experts, which considered context-grounded explanations highly informative for consistency and improving human-assessment. Table 3 summarizes results on 9 datasets synthesized based on humanitarian scope, and 14 real humanitarian datasets from HDX without sensitive data to evaluate false positives.
7 Contextual Sensitive Data: importance, practice, and consequences
With the emergence of public data sharing portals, and their adoption in critical areas such as the Humanitarian Data Exchange (HDX) for humanitarian aid coordination, it is key to protect sensitive datasets before they are automatically retrieved from the portal by malicious users or web crawlers for AI training datasets. Not doing so, induces significant risk to vulnerable populations. We reconsider and broaden the definition of what data is sensitive, and argue that sensitive data should be assessed contextually. This shift is critical as prior research for detecting and handling sensitive data typically focuses mainly on personal-level sensitive data. To broaden this practice, we present mechanisms for detecting sensitive data based on the principles of type contextualization and domain contextualization, which form a steppingstone towards this ambition.
On real-world humanitarian datasets, we show that more contextual methods lead to higher precision, hence less false positive alarms, while preserving coverage. These mechanisms can enable automation but also assist humans in data reviewing by providing context-grounded model explanations and concrete citations to relevant sensitivity rules in domain-specific documents. For successful adoption in practice, reasonable sensitivity instructions should be available in the form of 1) fine-grained semantic types of data that are deemed sensitive for type contextualization, and 2) reasonably specified assessment rules in domain-specific documents.
A potential negative consequence of automated data sensitivity detection is that critical datasets might be blocked that are time-sensitive and could be instrumental for solving societal problems, for example, in humanitarian domains. The mechanisms could be tailored around the desired precision/recall trade-off to reduce false alarms. At this moment, the methods are only developed for data sensitivity detection, and not for automated remedies. In practice, a report could be generated based on existing cases to instruct the user on how to remedy the sensitive data or potentially even automatically remedy this.
8 Conclusion
The rise of open data portals are key in facilitating the democratization of access to important data. This development, while positive, comes with risks of having sensitive data being used for harmful purposes by malicious actors as well as indirect leakage into language models which are fueled by any data on the web. Protecting sensitive data from getting published through such portals, by accident, is therefore key. In this paper, we first revisit the concept of sensitive data which is typically focused on personal-level data like personal-identifiable information (PII). Methods for detecting sensitive data are limited: they are too specific to cover all sensitive data, and too generic to only flag data that is actually sensitive. We have introduced a broader, contextual, perspective on sensitive data, and present a framework that formalizes contextual sensitive data from two angles. Experiments show that type-contextualization through an LLM-assisted mechanism that first detects the type of the data values (e.g. names) followed by a reflection on their sensitivity, significantly improves precision and recall over commercial tools and out-of-the-box LLM baselines. A case study with humanitarian datasets on the Humanitarian Data Exchange shows that the domain-contextualization through automated retrieval of rule-based heuristics from domain-specific documents effectively grounds an LLM’s assessment of data sensitivity. With our conceptualization of contextual sensitive data, we intend to encourage researchers and practitioners to consider a broader view on data sensitivity. We believe that our framework for detecting sensitive data provides a stepping stone in this important direction.
Acknowledgements
We are grateful for the collaboration with the Responsible Data team of the United Nations OCHA Humanitarian Data Centre. Specifically, we like to thank Alexandru Gartner, Godfrey Takavarasha, Jos Berens, Javier Téran, Metasebya Sahlu, Melanie Rabier, Nafissatou Pouye, Sarah Telford, Serban Teodorescu, and Shahrooz Badkoubei, for their invaluable contributions to this project. This work has been supported by funding from an AiNed grant from NWO (NGF.1607.22.045), a grant from SAP, and credit grants from OpenAI, Google and Cohere.

### v2

@@
 The emergence of open data portals necessitates more attention to protecting sensitive data before datasets get published and exchanged.
-While an abundance of methods for suppressing sensitive data exist, the conceptualization of sensitive data and methods to detect it, focus particularly on personal data that, if disclosed, may be harmful or violate privacy.
-We observe the need for refining and broadening our definitions of sensitive data, and argue that the sensitivity of data depends on its context.
-Based on this definition, we introduce two mechanisms for contextual sensitive data detection that consider the broader context of a dataset at hand.
-First, we introduce type contextualization, which first detects the semantic type of particular data values, then considers the overall context of the data values within the dataset or document.
-Second, we introduce domain contextualization which determines sensitivity of a given dataset in the broader context based on the retrieval of relevant rules from documents that specify data sensitivity (e.g., data topic and geographic origin).
-Experiments with these mechanisms, assisted by large language models (LLMs), confirm that: 1) type-contextualization significantly reduces the number of false positives for type-based sensitive data detection and reaches a recall of 94% compared to 63% with commercial tools, and 2) domain-contextualization leveraging sensitivity rule retrieval is effective for context-grounded sensitive data detection in non-standard data domains such as humanitarian datasets.
-Evaluation with humanitarian data experts also reveals that context-grounded LLM explanations provide useful guidance in manual data auditing processes, improving consistency.
-We open-source mechanisms and annotated datasets for contextual sensitive data detection at https://github.com/trl-lab/sensitive-data-detection.
+To do so effectively, we observe the need to refine and broaden our definitions of sensitive data, and argue that the sensitivity of data depends on its context.
+Following this definition, we introduce a contextual data sensitivity framework building on two core concepts: 1) type contextualization, which considers the type of the data values at hand within the overall context of the dataset or document to assess their true sensitivity, and 2) domain contextualization, which assesses the sensitivity of data values informed by domain-specific information external to the dataset, such as geographic origin of a dataset.
+Experiments instrumented with language models confirm that: 1) type-contextualization significantly reduces the number of false positives for type-based sensitive data detection and reaches a recall of 94% compared to 63% with commercial tools, and 2) domain-contextualization leveraging sensitivity rule retrieval effectively grounds sensitive data detection in relevant context in non-standard data domains.
+A case study with humanitarian data experts also illustrates that context-grounded explanations provide useful guidance in manual data auditing processes.
+We open-source the implementation of the mechanisms and annotated datasets at https://github.com/trl-lab/sensitive-data-detection.
+Towards Contextual Sensitive Data Detection
+Liang Telkamp1,2 and Madelon Hulsebos2thanks: Correspondence to madelon.hulsebos@cwi.nl 1University of Amsterdam, 2CWI
 1 Introduction
-Data sharing underpins reuse across sciences, enterprises, and governments through public and private portals [borgman2025data, worth2024datatransparency, zhang2024datamarketssurvey, wilkinson2016fair, brickley2019googledataset].
+Data sharing underpins data reuse across sciences, enterprises, and governments through public and private portals (Borgman and Groth, 2025; Worth et al., 2024; Zhang et al., 2024; Wilkinson et al., 2016; Brickley et al., 2019).
 With the rise of open data sharing portals, comes the need to protect sensitive data from undesired and harmful use.
-Sensitive data includes direct personal identifiers (e.g., names, emails [gdpr]) as well as indirect or contextual information that becomes a potential risk under certain conditions [unocha2025].
+Sensitive data includes direct personal identifiers (e.g., names, emails (European Parliament and Council of the European Union, 2016)) as well as indirect or contextual information that becomes a potential risk under certain conditions (OCHA, 2025).
 For example, facility locations or military logs may be sensitive depending on geopolitics or timing.
 This illustrates the complexity of sensitive data detection and the need for nuanced, context-dependent methods.
-Beyond operational and ethical concerns of publishing sensitive data, Large Language Models (LLMs) increase the risks of leakage: trained on web data, they may memorize and reproduce sensitive information [subramani2023piihuggingface, worth2024datatransparency, carlini2021extracting, lukas2023analyzing].
-Studies show sensitive data is often underspecified in documentation on open data portals like HuggingFace [yang2024navigatingdatasetdocumentationsai, croissant2024benjelloun, worth2024datatransparency], illustrating the importance of accurate methods and tools for (semi-)automated sensitive data protection, particularly in open data portals.
-Most work on sensitive data detection methods, however, focuses on detecting person-level data, particularly personal identifiable information (PII) [KUZINA2023119924, subramani2023piihuggingface, croissant2024benjelloun].
-For unstructured data, PII detection relies on named entity recognition [subramani2023piihuggingface]; similar approaches exist for tabular data [raman2001potter, Hulsebos2019Sherlock], which is the focus in this paper.
+Beyond operational and ethical concerns of publishing sensitive data, Large Language Models (LLMs) increase the risks of leakage: trained on web data, they may memorize and reproduce sensitive information (Subramani et al., 2023; Worth et al., 2024; Carlini et al., 2021; Lukas et al., 2023).
+Studies show sensitive data is often underspecified in documentation on open data portals like HuggingFace (Yang et al., 2024; Akhtar et al., 2024; Worth et al., 2024), illustrating the importance of accurate methods and tools for (semi-)automated sensitive data protection, particularly on open portals.
+Most work on sensitive data detection methods, however, focuses on detecting person-level data, particularly personal identifiable information (PII) (Kužina et al., 2023; Subramani et al., 2023; Akhtar et al., 2024).
+For unstructured data, PII detection relies on named entity recognition (Subramani et al., 2023); similar approaches exist for tabular data (Raman, 2001; Hulsebos et al., 2019), which is the focus in this paper.
 Yet sensitivity extends far beyond explicit PII.
-kober2023whatissensitive propose a taxonomy of always-, combination-, context-, and value-sensitive data.
+Kober et al.
+(2023) propose a taxonomy of always-, combination-, context-, and value-sensitive data.
 For instance, an address may be harmless in the context of a public organizational website but expose a private address in another document, causing naïve type-based detection yield false positives.
-Likewise, hospital geo-coordinates may be harmless in one region but sensitive in conflict zones [unocha2025], calling for contextualization beyond detecting personal-sensitive data.
+Likewise, hospital geo-coordinates may be harmless in one region but sensitive in conflict zones (OCHA, 2025), calling for contextualization beyond detecting personal-sensitive data.
 These limitations call for a more nuanced perspective and framework for assessing sensitive data.
-LLMs offer new potential for sensitivity detection by reasoning over structured data.
-Studies show they can interpret tables with carefully designed prompts [fang2024largelanguagemodelsllmstabular], achieving strong results in tasks such as text-to-SQL and table retrieval [fang2024largelanguagemodelsllmstabular, ji2025target].
-These capabilities suggest that LLMs can leverage schema, values, and external knowledge to interpret tabular dataset semantics [Hulsebos2019Sherlock].
-In this paper, we first revisit the concept of sensitive data, arguing for a more contextual perspective.
-We introduce two new LLM-assisted mechanisms for contextual sensitive data detection.
-First, we introduce the detect-then-reflect mechanism for type contextualization of data values within the entire document or tabular dataset to assess their sensitivity.
-Experiments with real-world datasets containing personal identifiable information show that this mechanism yields a strong improvement in precision while maintaining high recall.
-Second, the retrieve-then-detect mechanism for domain contextualization which grounds the assessment in domain-specific sensitivity specifications.
+In this paper, we revisit the concept of sensitive data detection, arguing for a more contextual perspective.
+We particularly focus on two aspects of contextual sensitive data detection.
+First, type contextualization which takes a more nuanced view beyond the type of values (e.g.
+addresses) by taking the broader context of the data into account.
+We introduce the technical “detect-then-reflect” mechanism which first detects the type of the values at hand, then reflects on their actual sensitivity based on contextual signals within the document or dataset.
+Experiments with real-world datasets containing personal identifiable information (PII) show that this mechanism, instrumented by language models, yields a strong improvement in precision while maintaining high recall.
+Second, we posit domain contextualization which addresses sensitivity that depends on external domain-specific factors, beyond merely static types such as PII.
+We introduce the “retrieve-then-detect” mechanism which first retrieves relevant domain-specific information, such as applicable data policies, to then ground the data sensitivity detection in using a language model.
 Evaluations illustrate the potential of this mechanism, achieving high recall and improved consistency for detecting non-personal sensitive data.
-Importantly, the context-grounded domain-specific explanations were deemed highly useful by experts in humanitarian data assessment in a case-study.
-2 Background
-Conceptualization of Sensitive Data
+Importantly, the context-grounded domain-specific explanations are deemed highly useful by experts in humanitarian data assessment in a case-study.
+2 Revisiting Sensitive Data Detection
 The notion of “sensitive data” is multifaceted, spanning legal, theoretical, and domain-specific views.
 Legal frameworks often equate sensitivity with personal identifiable information.
-The GDPR defines personal data broadly as “any information relating to an identified or identifiable natural person,” specifying sensitive categories like health, ethnicity, or politics requiring extra protection [gdpr].
-Similarly, HIPAA defines identifiers that make health information “Protected Health Information (PHI)” when linked to individuals [hipaa].
-These protocols provide a baseline: data that directly or indirectly identifies a person is sensitive.
-Yet static, type-level definitions are insufficient for modern governance [Quinn_Malgieri_2021], as they emphasize direct identifiers while overlooking non-personal but risky data such as geolocation data, community-level attributes, or institutional records.
-Sensitivity is often contextual: seemingly harmless data can become sensitive when combined, repurposed, or shared across contexts [malkin2022contextualintegrity].
-Theoretical perspectives deepen this nuance.
-Nissenbaum’s Contextual Integrity frames sensitivity as the appropriateness of information flows within social norms [nissenbaum2004privacy].
-Violations occur when flows breach expectations, not from the data type itself.
-Building on this, researchers advocate context-grounded definitions [malkin2022contextualintegrity, shvartzshnaider2019going].
-For instance, shopping data may be ordinary in retail but sensitive in behavioral profiling.
-Recent taxonomies further refine this view.
-kober2023whatissensitive distinguish: (1) always-sensitive data (e.g., direct identifiers); (2) combination-sensitive (sensitive when linked together); (3) context-sensitive (sensitivity by use); and (4) value-sensitive (sensitivity from specific values).
-This shows that not all personal is equally sensitive, and not all sensitive data is on individuals.
-In short, conceptualizations have shifted from static identifiers to context-dependent models, where frameworks like Contextual Integrity and multi-layer taxonomies emphasize context, use, and value.
-This motivates our focus: effective data protection must ensure coverage beyond personal-level information toward context-aware, domain-specific approaches.
-Methods for Detecting Sensitive Data
-Most automated sensitive data detection methods concentrate on detecting sensitive data of individuals.
-Widely adopted industry tools like Microsoft Presidio [presidio] and Google Cloud DLP [google_DLP] are largely rule-based, relying on regular expressions, keyword lists, and simple machine learning heuristics to recognize well-defined patterns like email addresses or phone numbers.
-Research has advanced towards more semantic understanding.
-Hulsebos2019Sherlock introduced Sherlock, a deep learning model that classifies types of table columns, including numerous PII, by learning from column-level representations.
-KUZINA2023119924 develop this further with CASSED, a BERT-based model that integrates column headers, data types, and sample values to classify columns into sensitive categories.
-These approaches outperform regular expressions and dictionary-based baselines, especially on noisy or inconsistently formatted data.
-Recently, Large Language Models (LLMs) have emerged as powerful tools for sensitive data detection, as particularly shown for PII detection outperforming both rule-based and traditional machine learning methods in accuracy and adaptability[shen2025enhancingdeidentificationpersonallyidentifiable, Yang2023Exploring].
-However, a key limitation remains: LLMs excel at identifying known PII types but struggle with assessing contextual sensitivity, such as domain-specific sensitive types, unless explicitly provided with the relevant operational context.
-We identify a significant gap in detecting sensitive data: methods focus on detecting personal sensitive data, overlooking non-personal data that is sensitive due to its context.
-Our work bridges this gap by developing a framework that addresses both personal and non-personal sensitive data detection through contextualization.
-3 Revisiting the Concept of Sensitive Data
-The main conceptual, legal, and methodological focus regarding sensitive data is on personal- and individual level data, including personal identifiable information (PII) which refers to data that can identify an individual [gdpr].
-Regulatory data protocols, such as GDPR [gdpr] and HIPAA [hipaa], specify standards for data sensitivity, hence provide concrete guidelines for data protection measures.
-However, we argue for the need of a more contextual view of sensitive data (Def.
-1) for developing new detection methods.
-Definition 1(Contextual Sensitive Data)
+The GDPR defines personal data broadly as “any information relating to an identified or identifiable natural person,” specifying sensitive categories like health, ethnicity, or politics requiring extra protection (European Parliament and Council of the European Union, 2016).
+Similarly, HIPAA defines identifiers that make health information “Protected Health Information (PHI)” when linked to individuals (Centers for Medicare & Medicaid Services, 1996).
+Yet, sensitivity is often contextual: seemingly harmless data can become sensitive when combined, repurposed, or shared across contexts (Malkin, 2022).Theoretical perspectives deepen this nuance.
+Nissenbaum’s Contextual Integrity frames sensitivity as the appropriateness of information flows within social norms (Nissenbaum, 2004).
+Building on this, researchers advocate context-grounded definitions (Malkin, 2022).Kober et al.
+(2023) distinguish: (1) always-sensitive data (e.g., personal identifiers); (2) sensitivity by linking data; (3) sensitivity by use; and (4) sensitivity from specific values.
+This echos that not all personal is equally sensitive, and not all sensitive data is on individuals.
+To detect sensitive data, of any kind, our framework builds on this conceptual notion of data sensitivity (Def.
+1): data protection methods must ensure coverage beyond personal- and static sensitivity definitions.
+Definition 1 (Contextual Sensitive Data).
 Data whose sensitivity depends on external factors and requires protection due to how, by whom, and in what context it can be misused.
-Advancing this definition, we posit two contextual aspects that matter in order to assess sensitivity of data: type contextualization and domain contextualization.
-Following the existence of generic standards and protocols, sensitive data detection methods often focus on detecting data types reflecting personal sensitive data (e.g., names, addresses, phone numbers, birth dates).
+Informed by this contextual definition of sensitive data, we formalize the problem of contextual sensitive data detection as follows:
+Problem Statement 1 (Contextual Sensitive Data Detection).
+Given an unordered or ordered sample of values $V={v_{1},v_{2},…,v_{N}}$ from document D, where each value $v_{i}$ represents a value, the set of internal context factors $C_{I}$ within $\mathcal{D}$ (e.g., surrounding values, structure), and external context factors $C_{E}$ outside $\mathcal{D}$ (e.g., specified rules, time/location), the objective of contextual sensitive data detection is to estimate the probability $P(V=sensitive|C_{E},C_{I})\in[0,1]$, and, based on this probability, classify the sensitivity of V into one of the predefined sensitivity levels $s_{1},s_{2},…,s_{M}$}, where each $s_{i}$ corresponds to a distinct sensitivity level (e.g., "sensitive", “medium sensitive”, "non-sensitive", etc.).
+3 Contextual Sensitive Data Detection
+Protecting sensitive data generally follows a three-step procedure: ➀ specifying sensitive data, ➁ detecting sensitive data, ➂ remedying data sensitivity.
+In this work, we focus on the second step, which has received little attention in research.
+Existing methods for detecting sensitive data suffer from two key limitations to address the contextual nature of sensitive data:
+  1. 1.
+    Insufficient specificity: sensitive data detection methods treat all values of types of interest as equally sensitive, while the subject of the value implies its sensitivity.
+This indistinctness leads to low precision hence high false positive rates.
+  2. 2.
+    Insufficient coverage: sensitive data detection methods typically focus on personal-level data, particularly PII, and do not generalize to domain-specific sensitive data.
+The lack of coverage leads to low recall of methods, hence leaving sensitive data unprotected.
+To address these challenges in contextual sensitive data detection, we posit two contextual aspects that matter in order to assess the contextual sensitivity of data: type contextualization and domain contextualization.
+We introduce two mechanisms to consider these contextual aspects of sensitive data: the detect-then-reflect mechanism for type contextualization and the retrieve-then-detect mechanism for domain contextualization.
+In what follows, we specifically focus on tabular data but these mechanisms extend to other modalities, such as free-form text documents.
+3.1 Type Contextualization
+Following existing standards and protocols European Parliament and Council of the European Union (2016); Centers for Medicare & Medicaid Services (1996), sensitive data detection methods often focus on detecting data types reflecting personal sensitive data (e.g., names, addresses) Google (2018).
 But the appearance of a name or address does not immediately make a data value sensitive.
@@
 We term this type contextualization: whether a data value corresponding to a sensitivity type of interest (e.g.
-name) truly is sensitive requires considering the context in which it appears in the entire dataset (e.g., the entire document or table).
-Besides a more fine-grained contextual notion of a data value its sensitivity, we also argue for a more coarse-grained contextual perspective.
-Nissenbaum introduced the theory of Contextual Integrity [nissenbaum2004privacy], which states that privacy risks arise when data is used outside its intended setting, particularly addressing personal privacy.
-This contextual perspective extends beyond personal privacy, in domains such as humanitarian aid, national security, and company-sensitive context.
-We pose that data protection should account for risks beyond identity exposure, for example, harming vulnerable populations or organizations.
-We refer to this broader perspective on data sensitivity as domain contextualization: sensitivity emerges from the dataset’s nature and potential misuse.
-4 Toward Contextual Sensitive Data Detection
-Protecting sensitive data generally follows a three-step procedure: ➀ specifying sensitive data, ➁ detecting sensitive data, ➂ remedying data sensitivity (Figure 1).
-In this work, we focus on the second step, which has received little attention in research.
-Existing methods for detecting sensitive data suffer from two key limitations to address the contextual nature of sensitive data:
-  1. 1.
-    Insufficient specificity: sensitive data detection methods treat all values of types of interest as equally sensitive, while the subject of the value that implies its sensitivity.
-The lack of specificity leads to low precision hence high false positive rates.
-  2. 2.
-    Insufficient coverage: sensitive data detection methods typically focus on personal-level data, particularly PII, and do not generalize to domain-specific sensitive data.
-This lack of coverage leads to low recall of methods, hence leaving sensitive data unprotected.
-Figure 1: Data protection process, starting with the specification of sensitivity, detection sensitive data in inputs accordingly, followed by remedying sensitive data by suppression methods.
-The contextualization framework introduced in Section 3 addresses these limitations.
-Type contextualization would improve specificity as potential sensitive values are only considered sensitive based on the broader context of the value.
-In turn, domain contextualization ensures higher coverage of sensitive data assessment by grounding the assessment in domain-specific context.
-In this paper, we introduce two concrete mechanisms to address these contextual aspects of sensitive data: the detect-then-reflect mechanism for type contextualization and the retrieve-then-detect mechanism for domain contextualization.
-In what follows, we specifically focus on tabular data but these mechanisms extend to other modalities, such as free-form text documents.
-(a) The detect-then-reflect mechanism for detecting sensitive table columns using type contextualization to increase the precision.
-This mechanism entails: ➀ detecting potentially sensitive columns based on their type, ➁ having an LLM reflect on the actual sensitivity of the column within the context of the entire table.
-(b) The retrieve-then-detect mechanism for detecting sensitive data that requires domain contextualization.
-This mechanism entails: ➀ retrieving relevant contextual information such as rule-based heuristics, ➁ using an LLM to detect sensitive data using its parametric knowledge and the retrieved expert heuristics.
-Figure 2: Mechanisms for detecting sensitive data in a contextual manner.
-4.1 Type contextualization: detect-then-reflect
-Existing methods typically classify sensitive data based on signals indicating a type of interest, e.g.
+name) truly is sensitive requires considering the context within the document or dataset in which it appears.
+Type contextualization improves specificity as potential sensitive values are only considered sensitive based on the broader context of the value.
+The detect-then-reflect mechanism
+Existing methods classify sensitive data based on signals indicating a type of interest, e.g.
 names, addresses, or phone numbers.
 If a value matches this type, it is flagged as sensitive.
-We introduce a two-step pipeline for detecting type-informed sensitive data based on a detect-then-reflect procedure (Figure 2(a)).
-In the initial detection step, potentially sensitive columns are identified by first classifying columns corresponding to types of interest that are specified upfront, e.g.
-email, phone number, based on column names and data values.
-This step prioritizes recall in order to avoid missing any columns that may carry data of the type that can potentially be sensitive.
-In the second step, reflection, the model assesses the columns in the context of the full table.
-We prompt an LLM with the table, including the column’s detected type.
-The LLM determines whether the column is truly sensitive, taking into account inter-column relationships and data across the entire table.
-This step is crucial for mitigating false positives from the detection stage by incorporating inter-column relationships and table semantics.
-4.2 Domain contextualization: retrieve-then-detect
-While type-based data sensitivity can often be inferred from structural context, the data sensitivity may depend on domain-specific and geopolitical factors that lie beyond the data itself.
-For such cases, requiring domain contextualization, we introduce the retrieve-then-detect mechanism (Figure 2(b)).
-This mechanism considers external contextual information, such as rule-based heuristics specified in data governance documents, in the classification process.
-Relevant documents are retrieved based on, for example, the origin of a dataset and included in the prompt alongside the table and column under analysis.
-The LLM leverages both its internal parametric knowledge and the retrieved external data sensitivity instructions to assess whether a column contains data that should be considered sensitive in the given operational context.
-Unlike type-informed detection, where there is a predefined list of entities (e.g.
-email, phone number) as relevant for for example PII detection, domain-contextual sensitive data has no fixed ontology.
-This mechanism establishes domain-specific column sensitivity based on the broader table context and retrieved contextual knowledge.
+Type contextualization employs a more nuanced assessment, for which we introduce the two-stage detect-then-reflect mechanism for detecting type-informed sensitive data.
+Algorithm 1 illustrates this staged procedure for binary sensitivity classification.
+Algorithm 1 Type Contextualization
+1:Values $V_{i}=\{v_{1},\dots,v_{N}\}$ from document $D$
+2:Internal context $C_{I}$, sensitive types $T_{\text{sens}}$, threshold $t$
+[…]
+4:if $t_{i}\in T_{\text{sens}}$ then
+5:  $p_{\text{sens}}\leftarrow\textsc{Estimate}$ $P(V_{i}=\text{sensitive}\mid t_{i},C_{I})$
+6:  $s_{i}\leftarrow(p_{\text{sens}}>t)$
+7:else
+8:  $s_{i}\leftarrow$ false
+9:end if
+In the initial detection step, the potential sensitivity of a set of values is estimated by first classifying the values ($V_{i}$) corresponding to sensitive types of interest specified upfront ($T_{\text{sens}}$).
+This step prioritizes recall by capturing all values that may carry data of the sensitive type.
+The purpose of the reflection step is to estimate the actual sensitivity of the values ($p_{\text{sens}}$) given their detected sensitive type ($t_{i}$) and internal context of the full document ($C_{I}$), to finally turn into a discrete sensitivity label ($s_{i}$).
+This step is crucial for mitigating false positives from the type detection stage by incorporating inter-value relationships and document semantics, to filter out values of sensitive types that are not actually sensitive.
+Tabular Data case.
+Figure 1 illustrates the detect-then-reflect mechanism for type contextualization on a concrete example with tabular data.
+Here, table columns are classified to correspond to any sensitive types such as address and phone number based on the column name and data values.
+For columns of sensitive types, the reflection step assesses whether the column is sensitive in context of the full table context, i.e., the table name as well as column names and values of surrounding columns.
+Figure 1: Type contextualization for contextually detecting sensitive table columns using the detect-then-reflect mechanism to increase precision.
+This mechanism entails: ➀ detecting potentially sensitive columns based on their (semantic) type, ➁ reflecting on the actual sensitivity of the column within the internal context of the entire table.
+3.2 Domain Contextualization
+Besides a more fine-grained contextual notion of the sensitivity of certain values, we also argue for a more coarse-grained contextual perspective.
+Following our notion of contextual sensitive data (Def.
+1), we pose that data protection should account for risks beyond identity exposure, for example, harming vulnerable populations or organizations hence consider the data domain, from humanitarian aid to national security.
+We refer to this broader perspective on data sensitivity as domain contextualization: sensitivity emerges from the dataset’s nature and potential misuse, which can arise from external factors.
+As such, domain contextualization aims at higher detection coverage by grounding the sensitivity assessment in domain-specific context, beyond generic and static predefined types.
+The retrieve-then-detect mechanism
+While type-based data sensitivity can often be inferred from internal document context, domain-specific sensitivity typically depends on contextual factors external to the data itself.
+For domain contextualization, we introduce the retrieve-then-detect mechanism (Alg.
+2).
+Algorithm 2 Domain Contextualization
+1:Values $V_{i}=\{v_{1},\dots,v_{N}\}$ from document $D$
+2:External context $C_{E}$, threshold $t$
+[…]
+[…]
+[…]
+Unlike type-informed detection, based on a predefined set of sensitive types, domain-specific sensitive data has no fixed ontology.
+Instead, the retrieve-then-detect mechanism considers external contextual information, such as rule-based heuristics specified in data governance documents or geopolitical news items, in the detection process.
+In the retrieval stage, external context is selected from a corpus of domain-specific documents or other external data sources ($C_{E}$) based on the relevance to the potential sensitive values at hand ($V_{i}$).
+In the detection stage, the sensitivity of the values at hand ($s_{i}$) is assessed given the retrieved relevant external context ($C_{i}^{\text{relevant}}$).
+Tabular Data case.
+Figure 2 illustrates the retrieve-then-detect mechanism for tabular data.
+For a given column (or table), the retrieval component fetches information, such as data policy rules or geopolitical information, from external documents or other data sources that are relevant for the given input table.
+The detection step then assesses whether a table (column), or combination of columns, contains data that should be considered sensitive given the retrieved relevant context.
+Figure 2: The retrieve-then-detect mechanism for detecting sensitive data that requires domain contextualization.
+This mechanism entails: ➀ retrieving relevant contextual information such as data policy rules, ➁ detecting sensitive data based on the table and the retrieved relevant context.
 4 Experimental Setup
-Here, we outline the implementation details, the LLMs used, baselines, datasets, and metrics to evaluate our proposed framework for contextually sensitive data detection.
+Here, we outline the method implementation details, baselines, datasets, and metrics to evaluate our proposed framework for contextually sensitive data detection.
 Code and datasets can be found at https://github.com/trl-lab/sensitive-data-detection.
-5.1 Method implementation
-5.1.1 Type-contextualization
-In the detection step, given the well-defined nature of PII types, we parse the problem as a multiclass classification task over a taxonomy of personal identifiable types (e.g., email_address, generic_identifier) derived from GDPR [gdpr] and UN OCHA guidelines [OCHA2021Guidelines] (see the code repository for a complete list).
-The LLM is prompted with a column name and five values, to return only the most appropriate type or None.
+4.1 Implementation of the Mechanisms
+The core components of the mechanisms are type detection, reflection, and retrieval. Further details of the implementation of these components per mechanism are provided below.
+Type-contextualization
+In the detection step, given the well-defined nature of PII types, we parse the problem as a multiclass classification task over a taxonomy of personal identifiable types (e.g., email_address, generic_identifier) derived from GDPR European Parliament and Council of the European Union (2016) and UN OCHA guidelines UN OCHA Centre for Humanitarian Data (2021) (see the code repository for a complete list).
+We devise a language model to classify the type of a table column (or return None) Kayali et al.
+(2024); Shen et al.
+(2025), by prompting it with the column name and five values.
 We consider zero-shot and fine-tuned approaches.
 The zero-shot method leverages the model’s inherent ability to follow instructions, while the fine-tuned approach enhances performance by training the models on a synthetic dataset of 1,000 labeled columns.
-Fine-tuning is done using Low-Rank Adaptation [LORA21_HU] for parameter efficiency.
-The models are optimized with AdamW for a single epoch using an 80/20 train-validation split.
+Fine-tuning is done using Low-Rank Adaptation Hu et al.
+(2021) for parameter efficiency.
+The models are trained for a single epoch using an 80/20 train-validation split.
 In the reflection phase, the LLM re-evaluates each candidate column corresponding to types of interest within the context of the entire table.
-We provide the LLM with a markdown-formatted representation of the full table, including headers and five sample rows per column, along with the specific type detected for the target column and, if applicable, the entities of other columns which are detected as PII entities.
+We provide the LLM with a markdown-formatted representation of the full table, including headers and five sample rows per column, along with the specific type detected for the target column and, if applicable, the types of other columns which are detected as PII types.
 The LLM is then instructed to assign one of three contextual sensitivity levels: non_sensitive, referring to information that cannot identify a person (e.g., aggregate data, organization names); moderate_sensitive, referring to data that could potentially identify a person when combined with other attributes (e.g., demographics, partial information); and high_sensitive, referring to data that definitively identifies a person (e.g., full name, email, national ID).
-5.1.2 Domain-contextualization
-For the retrieval step, we aim to identify relevant heuristics indicating sensitivity rules from domain-specific data governance documents in order to integrate this external domain knowledge in the LLM’s context.
-In our experiments we focus on a case study on the Humanitarian Data Exchange (HDX) [hdx], hence on humanitarian datasets.
+Domain-contextualization
+Given the domain-specific nature of this mechanism, we conduct a case study with humanitarian data experts from the UN OCHA Humanitarian Data Centre using humanitarian crisis datasets from the Humanitarian Data Exchange (HDX) United Nations (2014).
+For the retrieval step, we aim to identify relevant heuristics indicating sensitivity rules from domain-specific data governance documents.
 The system retrieves relevant policy fragments from Information Sharing Protocols (ISPs) based on the dataset’s geographical context (e.g., country).
-We use GPT-4o-mini to extract the specific sensitivity rules from these documents (e.g., data types for low, moderate, high, severe sensitivity), which we store in a dictionary.
-If no country-specific ISP is identified, a default fallback template based on cross-country consensus is used.
+We use GPT-4o-mini to extract and index the specific sensitivity rules from these documents.
+If no region-specific ISP is identified, a default ISP template based on cross-country consensus is used.
 In the detection phase, the LLM is provided with a curated prompt containing table context (headers and five representative rows), relevant ISP extracts containing explicit instructions to classify each column’s sensitivity according to the ISP guidelines while justifying its decision by directly citing the relevant policy passages.
@@
 4.2 Model Selection
-We evaluate closed-source and open-source models, to understand trade-offs between these options.
-From closed-source options, we select the GPT-4o-mini [openai2024gpt4ocard] model for its good performance and cost trade-off.
-We focus on open-source LLMs that enable local deployment, for which we select the Gemma family (2 and 3) [gemma_2025, Gemma2Paper] and Qwen3 in 8B and 14B sizes [yang2025qwen3technicalreport].
+We evaluate closed-source and open-source models to understand trade-offs between these options, if any.
+While one of the goals of this work is to prevent sensitive data from leaking into proprietary LLMs, these LLMs are typically available under privacy-preserving protocols at some cost which can then provide a viable low-maintenance solution.
+From closed-source options, we select the GPT-4o-mini (OpenAI et al., 2024) model for its good performance and cost trade-off.
+We focus on open-source LLMs that enable local deployment, for which we select the Gemma family (2 and 3) (Team, 2025; Team et al., 2024) and Qwen3 in 8B and 14B sizes (Yang et al., 2025).
 Qwen3 14B was selected given the strong performance of its predecessor Qwen2.5’s exceptional performance on benchmarks such as MMLU Pro  [^1] https://huggingface.co/spaces/TIGER-Lab/MMLU-Pro, demonstrating strong reasoning capabilities and robust general understanding.
 Furthermore, Gemma2 9B and Gemma3 12B were included to strike a balance between computational efficiency and performance, and their support for quantized deployment.
-We also include versions of Qwen3 8B and Gemma2 9B that are fFgine-tuned on a synthetic dataset of 1K table columns hand-labeled with 27 PII types, indicated with “FT” suffix.
+We also include versions of Qwen3 8B and Gemma2 9B that are fine-tuned on a synthetic dataset of 1K table columns hand-labeled with 27 PII types, indicated with “FT” suffix.
 4.3 Baselines
 We compare our methods with two common baselines for sensitive data detection.
-We include Google Cloud DLP [google_DLP], a rule-based system that combines built-in infoTypes, predefined detectors for common PII entities, which are associated custom-defined dictionaries and regular expressions.
+We include Google Cloud DLP Google Cloud (2018), a rule-based system that combines built-in infoTypes, predefined detectors for common Personal Identifiable Information (PII) types, which are associated custom-defined dictionaries and regular expressions.
 For our experiments, we configure the minimum likelihood threshold to “Likely”.
@@
 To evaluate the contextual sensitive data detection mechanisms we use a combination of synthetic and real-world datasets, to maintain ground-truth validity for sensitive data detection in generic context and the domain-specific context of humanitarian datasets.
-  • Type contextualization with sensitive PII: We extracted 66 anonymized tables (2,061 columns) from the GitTables corpus [hulsebos2023gittables] representing a variety of real-world database schemas.
-Each column was manually annotated with (i) its PII entity type and (ii) its contextual sensitivity level (non-sensitive, moderate or highly sensitive).
-  • Domain contextualization with humanitarian tables: To evaluate the retrieve-then-detect mechanism, we generated 9 synthetic tables with sensitive data closely resembling typical humanitarian data (e.g., for displacement tracking, public health monitoring) for different countries.
-Each table contains columns that could be potentially sensitive within country-specific Information Sharing Protocols (ISPs).
-Humanitarian data quality experts labeled each column with a sensitivity level (low, moderate, high, severe).
-We also add 14 real tables from the Humanitarian Data Exchange [^2] https://data.humdata.org/ that do not contain any sensitive data, to evaluate the false positive rate.
+  • Type contextualization with sensitive PII: We extracted 66 anonymized tables, with in total 2,061 columns, from the GitTables corpus Hulsebos et al.
+(2023) representing a variety of real-world table schemas.
+Each column was manually annotated with (i) its PII type and (ii) its contextual sensitivity level (non-sensitive, moderate, or highly sensitive).
+  • Domain contextualization with humanitarian data: To evaluate the retrieve-then-detect mechanism for domain-specific sensitive data, we conduct a case study with humanitarian experts from the UN OCHA Humanitarian Data Centre.
+We collect 23 tabular datasets, with in total 326 columns, resembling typical humanitarian datasets (e.g., for displacement tracking, public health monitoring) for different regions.
+With rich input from humanitarian experts to ensure realistic properties, we synthesize 9 tabular datasets with one or more actual contextually sensitive data in line with region-specific Information Sharing Protocols (ISPs).
+Two humanitarian data experts annotated columns (100 in total, of which 70 contain sensitive data) with a sensitivity level (low, moderate, high, severe).
+The other 14 tabular datasets are similar but real tabular datasets from the UN’s Humanitarian Data Exchange [^2] https://data.humdata.org/ and have been manually curated as challenging datasets without actual sensitive data, to evaluate the false positive rate.
+Evaluation granularity
+For evaluation, sensitivity levels are binarized: low is treated as non-sensitive, while moderate, high, and severe are grouped as sensitive.
+This aligns with operational practice in humanitarian contexts where any moderate or higher sensitivity triggers safeguards, consistent with UN OCHA guidance that such data warrants protective measures (OCHA, 2025).
+For domain contextualization, we assess performance on table-level to align with the ISP’s coarse-grained data policy instructions, despite having column-level annotations.
+Tables are labeled sensitive if one of its columns is annotated as sensitive (moderate, high or severe).
 4.5 Metrics
-We evaluate along three axes: PII detection, contextual reflection for PII, and non-PII sensitivity detection.
-To account for class imbalance, we report both weighted (support-weighted) and macro averages.
-Weighted averages reflect overall performance across examples, while macro averages treat each class equally and reduce dominance by frequent classes.
-Unless otherwise noted, we focus on weighted averages.
-Performance is assessed at both the column and table level.
-At the column level (type contextualization), PII detection is treated as a multiclass task over 27 PII types (plus a “None” category), while sensitivity reflection reduces to a binary decision between sensitive (moderate/high) and non-sensitive (low).
-At the table level (domain contextualization), a table is classified as sensitive if it contains any moderate/high/severe sensitive information, resulting in a binary task (sensitive vs.
+We evaluate accuracy along three axes: PII detection, contextual reflection for PII, and non-PII sensitivity detection.
+To account for class imbalance, we report (support-)weighted averages.
+Performance is measured at column and table level.
+For type contextualization we evaluate column-level accuracy, PII detection is treated as a multiclass task over 27 PII types (plus a “None” category), while sensitivity reflection reduces to a binary decision between sensitive (moderate/high) and non-sensitive (low).
+For domain contextualization we evaluate table-level accuracy: a table is classified as sensitive if it contains any moderate/high/severe sensitive information, resulting in a binary task (sensitive vs.
 not).
 Precision, recall, and F1 are used as evaluation metrics.
-For labeling, sensitivity levels are binarized: low is treated as non-sensitive, while moderate, high, and severe are grouped as sensitive.
-This aligns with operational practice where any moderate or higher sensitivity triggers safeguards, consistent with UN OCHA guidance that such data warrants protective measures [unocha2025].
-6 Results of contextual sensitive data detection
+Besides accuracy metrics, we report the average runtime per column to assess deployment considerations.
+5 Results
 We present a comprehensive evaluation of our proposed framework for contextually sensitive data detection in tabular datasets, through type- and domain-contextualization.
-Type Contextualization experiments instrumented by the detect-then-reflect mechanism are focused on personal-identifiable information (PII) and Domain Contextualization instrumented by the retrieve-then-reflect mechanism which focus on less standard data semantics through a humanitarian case-study.
+Type Contextualization experiments instrumented by the detect-then-reflect mechanism are focused on personal-identifiable information (PII) and Domain Contextualization instrumented by the retrieve-then-detect mechanism which focuses on domain-specific sensitivity through a case-study with humanitarian data experts.
 5.1 Type contextualization improves specificity.
@@
 We observe that reflection consistently improves precision and F1.
-System / Model Without Reflection With Reflection
+system/model w/o reflection w/ reflection
 prec. rec. F1 prec. rec. F1
-Google DLP 0.531 0.628 0.576
-Presidio 0.520 0.618 0.565
-GPT-4o-mini 0.856 0.639 0.732 0.938 0.632 0.755
-Gemma 2 9B 0.740 0.819 0.778 0.800 0.792 0.796
-Gemma 3 12B 0.487 0.941 0.642 0.753 0.806 0.779
-Qwen3 8B 0.742 0.868 0.800 0.749 0.868 0.804
-Qwen3 14B 0.565 0.972 0.714 0.732 0.941 0.824
-Qwen3 8B FT $\rightarrow$ GPT-4o-mini 0.902 0.861 0.881
-Without reflection, LLMs mark any detected PII-type column as sensitive, leading to high recall but lower precision; Gemma 3 12B and Qwen3 14B reach a recall of 0.94 and 0.97 but precision of only 0.49 and 0.57, respectively.
+Google DLP 0.53 0.63 0.58
+Presidio 0.52 0.62 0.57
+GPT-4o-mini 0.86 0.64 0.73 0.94 0.63 0.76
+Gemma 2 9B 0.74 0.82 0.78 0.80 0.79 0.80
+Gemma 3 12B 0.49 0.94 0.64 0.75 0.81 0.78
+Qwen3 8B 0.74 0.87 0.80 0.75 0.87 0.80
+Qwen3 14B 0.57 0.97 0.71 0.73 0.94 0.82
+Qwen3 8B FT -> GPT-4o-mini 0.90 0.86 0.88
+Without reflection, LLMs mark any detected PII-type column as sensitive, leading to high recall but lower precision; Gemma 3 12B and Qwen3 14B reach a recall of 0.94 and 0.97 but precision of only 0.49 and 0.57 respectively (Table 1).
 GPT-4o-mini achieves higher precision, 0.86, at lower recall of 0.64.
@@
 Comparing the detect-then-reflect mechanism powered by open-source LLMs with commonly used pattern-matching tools DLP and Presidio, we observe a significant improvement in both precision and recall, addressing the key limitations of existing tools for deployment in practice.
-Overall, the detect-then-reflect mechanism corrects the naive assumption that all columns with potentially sensitive types are sensitive, boosting precision while maintaining recall.
-6.1.1 Ablation: reflection without detection.
+Overall, the detect-then-reflect mechanism corrects the naive assumption that all columns of sensitive types are indeed always sensitive, boosting precision while maintaining recall.
+Ablation: reflection without detection.
 To investigate the value of the type-detection step, we also task models with identifying sensitive columns using directly the full table context without first detecting the presence of columns of specific types (Table 1).
@@
 These results indicate that LLMs can reason about sensitivity from the table alone, but the full detect-then-reflect mechanism yields preferable trade-off in precision and recall.
-Table 2: Reflection-only performance without prior type detection step, for identifying contextually sensitive columns. Evaluated across models using precision, recall, and F1 scores.
+Table 2: Reflection-only performance without prior type detection, for identifying contextually sensitive columns. Evaluated across models using precision, recall, and F1 scores.
 model precision recall F1
-GPT-4o-mini 0.964 0.562 0.711
-Gemma 2 9B 0.878 0.750 0.809
-Qwen3 8B 0.735 0.781 0.758
-Qwen3 14B 0.761 0.885 0.819
+GPT-4o-mini 0.96 0.56 0.71
+Gemma 2 9B 0.88 0.75 0.81
+Qwen3 8B 0.74 0.78 0.76
+Qwen3 14B 0.76 0.89 0.82
 5.2 Domain contextualization provides coverage and context-grounded explainability.
 Shifting towards contextual sensitive data which cannot be assessed through a fixed set of types, we evaluate the retrieve-then-detect mechanism on a combined set of synthetic and real humanitarian tables from the HDX platform.
-Without external domain knowledge, i.e.
-sensitivity heuristics obtained through context retrieval, all models exhibit a highly conservative bias, achieving perfect recall but all with precision lower than 0.56 (Table 3).
-Integrating retrieved domain-specific context, in this case we leverage information sharing protocols (ISP’s) for humanitarian data sharing, into the prompt consistently improves precision.
+Without external domain knowledge, all models exhibit a highly conservative bias, achieving perfect recall but all with precision lower than 0.56 (Table 3; baseline).
+Integrating retrieved domain-specific context, i.e., sensitivity guidelines obtained from information sharing protocols (ISP’s) for humanitarian data sharing, into the input for reflection by the language model, consistently improves precision.
 GPT-4o-mini’s precision increases from 0.47 to 0.69 (F1=0.82), and Qwen3 14B’s from 0.56 to 0.64 (F1=0.78), both maintaining perfect recall (1).
+Table 3 summarizes the results on the 23 datasets humanitarian datasets.
 Table 3: Table-level classification performance with and without context-grounding in humanitarian data sensitivity rules. The retrieved rules improve precision while maintaining high recall.
-System / Model No Domain Knowledge With Domain Knowledge
+system/model w/o domain knowledge w/ domain knowledge
 prec. rec. F1 prec. rec. F1
-All-tables-sensitive (baseline) 0.375 1.000 0.545
+Baseline: all sensitive 0.375 1.000 0.545
 GPT-4o-mini 0.474 1.000 0.643 0.692 1.000 0.818
@@
 Qwen3 14B 0.562 1.000 0.720 0.643 1.000 0.783
-Qualitative analysis reveals that the retrieval mechanism enables models to produce interpretable justifications, thereby enhancing the transparency and auditability of the sensitivity assessments.
-The below excerpt shows a context-grounded LLM assessment on a table titled “Reported location on NFIs suppliers” with a description of the data at hand (1), along with the LLM’s sensitivity assessment (2), and explanation grounded in retrieved ISP rules (3).
-1. The column ‘‘nfi_supplier_loc- value -from_Iran’’ indicates the origin of NFIs suppliers.
-2. This information, even with country-level disaggregation, doesn’t reveal personal or sensitive data. It relates to the location of suppliers, which is generally not considered sensitive.
+Qualitative analysis surfaces that the context retrieval mechanism enables models to produce interpretable and consistent justifications, thereby enhancing the transparency and auditability of the sensitivity assessments.
+The model outputs containing sensitivity classifications and explanations were reviewed by two humanitarian data quality experts, who considered the context-grounded explanations accurate, and highly informative for guiding and standardizing human-assessment.
+The below excerpt shows such context-grounded LLM assessment on a table titled “Reported location on NFIs suppliers” with (1) a description of the data at hand, along with (2) the LLM’s sensitivity assessment, and (3) explanation grounded in retrieved ISP rules.
+1.
+The column ‘‘nfi_supplier_loc-value-from_Iran’’ indicates the origin of NFIs suppliers.
+2.
+This information, even with country-level disaggregation, doesn’t reveal personal or sensitive data.
+It relates to the location of suppliers, which is generally not considered sensitive.
 3.
 It aligns with the ISP’s guideline on ‘‘Facility data (e.g., health, education, water points) at national or regional level, unless explicitly restricted by clusters.’’ Since this column is about locations of suppliers and not about individuals or specific operations, it is categorized as non-sensitive.
-The model outputs containing sensitivity classifications and explanations were also reviewed by two humanitarian data quality experts, which considered context-grounded explanations highly informative for consistency and improving human-assessment.
-Table 3 summarizes results on 9 datasets synthesized based on humanitarian scope, and 14 real humanitarian datasets from HDX without sensitive data to evaluate false positives.
-7 Contextual Sensitive Data: importance, practice, and consequences
-With the emergence of public data sharing portals, and their adoption in critical areas such as the Humanitarian Data Exchange (HDX) for humanitarian aid coordination, it is key to protect sensitive datasets before they are automatically retrieved from the portal by malicious users or web crawlers for AI training datasets.
-Not doing so, induces significant risk to vulnerable populations.
-We reconsider and broaden the definition of what data is sensitive, and argue that sensitive data should be assessed contextually.
-This shift is critical as prior research for detecting and handling sensitive data typically focuses mainly on personal-level sensitive data.
-To broaden this practice, we present mechanisms for detecting sensitive data based on the principles of type contextualization and domain contextualization, which form a steppingstone towards this ambition.
-On real-world humanitarian datasets, we show that more contextual methods lead to higher precision, hence less false positive alarms, while preserving coverage.
-These mechanisms can enable automation but also assist humans in data reviewing by providing context-grounded model explanations and concrete citations to relevant sensitivity rules in domain-specific documents.
-For successful adoption in practice, reasonable sensitivity instructions should be available in the form of 1) fine-grained semantic types of data that are deemed sensitive for type contextualization, and 2) reasonably specified assessment rules in domain-specific documents.
+5.3 Error analysis
+Type contextualization
+On data with PII, we particularly observed misclassification of columns of generic types, as well as explicit personal types.
+For example, the “GENERIC_ID” type is well detected, but classified as non-sensitive by most LLMs causing very low recall.
+For types more explicitly relating to personal information (like names), columns are often classified as sensitive in reflection, while not always annotated as such.
+This highlights the potential for more specific context- and type-specific instructions during reflection.
+Domain contextualization
+In our experiments with humanitarian data, we observed mainly false positives for LLMs, which generally appear cautious with humanitarian data.
+On these false positives, analysis surfaces that the Qwen model has most confusion around borderline cases that are annotated as non-sensitive but classified as medium-sensitive, while gpt-4o-mini deems most of these tables highly sensitive (75%).
+This illustrates the need for more specific instructions for borderline cases.
+5.4 Latency of contextual sensitive data detection.
+Table 4 gives an overview of wall-clock latency measurements of the end-to-end pipeline averaged over all columns, as well as cost indications and memory measurements for local models.
+Despite the more elaborate mechanisms for contextualized sensitive data detection, the latency delta between the contextual sensitive LM-powered pipelines versus Google DLP (which only conducts type-based sensitivity detection) might be acceptable given the significant accuracy gain that the introduced mechanisms provide.
+If efficiency is key, smaller models like Qwen3 8B can yield an favorable trade-off, whereas larger or more costly models for more sensitive scenarios can be feasible for smaller-scale evaluations.
+Table 4: Comparison of pipeline latency per column and resource cost across different methods and models.
+tool/model total pipeline (s/col) cost or GPU memory
+Google DLP [^3] For Google DLP, we only measure type-based sensitivity detection as it is not suitable for domain-specific sensitivity. 0.16 subscription
+GPT-4o-mini 1.20 \$0.05 per table
+Gemma 2 9B 2.89 7.8 GB
+Gemma 3 12B 2.94 8.0 GB
+Qwen3 8B 0.46 8.0 GB
+Qwen3 14B 1.62 11.4 GB
+6 Related Work
+Existing methods for detecting sensitive data concentrate on detecting PII types, or otherwise static sensitive types.
+To this end, widely adopted industry tools like Microsoft Presidio (Microsoft, 2018) and Google Cloud DLP (Google Cloud, 2018) detect correspondences between values and structured types, such as email addresses and phone numbers, based on rule-based heuristics and syntactic pattern matching.
+Recent research has expanded the type detection scope by utilizing large neural models to learn mappings between high-dimensional semantic representations of values, e.g.
+table columns, and their semantic types Hulsebos et al.
+(2019); Kužina et al.
+(2023).
+Language models have also emerged as powerful tools for sensitive data detection, as particularly shown for PII detection, outperforming both rule-based and traditional machine learning methods in accuracy and adaptability (Shen et al., 2025; Yang et al., 2023).
+However, while language models excel at identifying known PII types, they struggle with assessing contextual sensitivity, such as domain-specific sensitive types, unless explicitly provided with the relevant operational context.
+In this work, we address this gap: our contextual sensitive data detection framework addresses both personal and non-personal sensitive data by considering internal and external context factors.
+7 Conclusion
+With the rise of open data portals to democratize data access, comes the risk of exposing sensitive data to malicious actors or language models fueled by public data.
+Protecting sensitive data is therefore key, but methods for this are limited: they are too specific to capture all sensitive data, and too generic to only suppress data that is actually sensitive.
+We argue for contextualizing sensitive data detection along two dimensions.
+First, type contextualization, which considers fine-grained aspects beyond static sensitivity types.
+Our detect-then-reflect mechanism shows that first detecting if data values belong to a sensitive type, then reflecting on their true sensitivity, increases recall over existing tools with up to 49%.
+Second, we introduce domain contextualization which considers domain-specific factors external to the data in order to assess its sensitivity.
+A case study with humanitarian datasets highlights that contextualizing the sensitivity assessment in relevant domain-specific rules through a retrieval stage, extends the type coverage and enables domain-grounded explanations.
+We believe that our work provides a stepping stone in this important direction.
+Limitations and Risks
+Evaluation coverage
+In our humanitarian case study, we show that domain contextualization leads to higher precision while preserving coverage, and also provides useful grounded explanations to data quality reviewers.
+The limited availability of labeled data in this case study, might affect generalization of the insights for domain-contextualized sensitive data detection.
+Moreover, while the conceptual framework and mechanisms are modality agnostic (Section 3), the lack of datasets with labeled sensitivity and the data setting of the humanitarian case study have resulted in evaluations for tabular data only, for which we constructed and publish hand-labeled datasets.
+We call for more diverse labeled datasets, to advance this important problem, and evaluate the framework and mechanisms for different modalities, languages, and domains.
+Modality coverage
+Despite the focus of our implementation and evaluation on tabular data.
+The proposed mechanisms easily extend to other modalities, such as free-form text data: 1) type-informed sensitivity in free-form text can be implemented efficiently by concatenating a decomposition component with a more lightweight learned type predictor Hulsebos et al.
+(2019) and performing reflection in batches, 2) for domain-contextualization of free-form text, a hierarchical reflection procedure could be efficient.
+For example, starting with assessing high-level segments of the text (document-level, paragraph-level, sentence-level, etc Lin et al.
+(2025)), and incrementally narrowing down.
+Sensitivity specification dependency
+The presented mechanisms for detecting sensitive data can enable automation but also assist humans in data reviewing by providing context-grounded explanations with explicit citations to relevant sensitivity rules.
+For successful adoption in practice, however, the mechanisms depend on having access to reasonable sensitivity instructions in the form of 1) fine-grained semantic types of data that are deemed sensitive for type contextualization, and 2) reasonably specified assessment rules in domain-specific documents or (web) data sources.
+Data access delay and further automation
 A potential negative consequence of automated data sensitivity detection is that critical datasets might be blocked that are time-sensitive and could be instrumental for solving societal problems, for example, in humanitarian domains.
 The mechanisms could be tailored around the desired precision/recall trade-off to reduce false alarms.
-At this moment, the methods are only developed for data sensitivity detection, and not for automated remedies.
-In practice, a report could be generated based on existing cases to instruct the user on how to remedy the sensitive data or potentially even automatically remedy this.
-8 Conclusion
-The rise of open data portals are key in facilitating the democratization of access to important data.
-This development, while positive, comes with risks of having sensitive data being used for harmful purposes by malicious actors as well as indirect leakage into language models which are fueled by any data on the web.
-Protecting sensitive data from getting published through such portals, by accident, is therefore key.
-In this paper, we first revisit the concept of sensitive data which is typically focused on personal-level data like personal-identifiable information (PII).
-Methods for detecting sensitive data are limited: they are too specific to cover all sensitive data, and too generic to only flag data that is actually sensitive.
-We have introduced a broader, contextual, perspective on sensitive data, and present a framework that formalizes contextual sensitive data from two angles.
-Experiments show that type-contextualization through an LLM-assisted mechanism that first detects the type of the data values (e.g.
-names) followed by a reflection on their sensitivity, significantly improves precision and recall over commercial tools and out-of-the-box LLM baselines.
-A case study with humanitarian datasets on the Humanitarian Data Exchange shows that the domain-contextualization through automated retrieval of rule-based heuristics from domain-specific documents effectively grounds an LLM’s assessment of data sensitivity.
-With our conceptualization of contextual sensitive data, we intend to encourage researchers and practitioners to consider a broader view on data sensitivity.
-We believe that our framework for detecting sensitive data provides a stepping stone in this important direction.
+In this work, we only present mechanisms for detecting contextual data sensitivity, and not for automated remedies.
+In practice, a report could be generated based on existing cases to instruct the user on how to remedy the sensitive data, or potentially even automatically remedy this, which is also an important direction for future research.
 Acknowledgements
 We are grateful for the collaboration with the Responsible Data team of the United Nations OCHA Humanitarian Data Centre.
-Specifically, we like to thank Alexandru Gartner, Godfrey Takavarasha, Jos Berens, Javier Téran, Metasebya Sahlu, Melanie Rabier, Nafissatou Pouye, Sarah Telford, Serban Teodorescu, and Shahrooz Badkoubei, for their invaluable contributions to this project.
-This work has been supported by funding from an AiNed grant from NWO (NGF.1607.22.045), a grant from SAP, and credit grants from OpenAI, Google and Cohere.
+Specifically, we thank Alexandru Gartner, Godfrey Takavarasha, Jos Berens, Javier Téran, Metasebya Sahlu, Melanie Rabier, Nafissatou Pouye, Sarah Telford, Serban Teodorescu, and Shahrooz Badkoubei, for their invaluable contributions to this project.
+This work has been supported by funding from an AiNed grant from NWO (NGF.1607.22.045) and credit grants from OpenAI.